I am still struggling with this topic. How would I create a quadratic equation with given solutions 9 and 16?
Found 2 solutions by solver91311, problemsolver:
Answer by solver91311(24713)



If is a zero of a polynomial function, , then must be a factor of the polynomial.

Hence, if your zeros are 9 and 16, your quadratic polynomial must be comprised of the factors , , and a real number constant .

John


Answer by problemsolver(1)

Using the principle of zero; you set both solutions of x to equal zero and work backwards.
x=9 x=16
x-9=0 x-16=0
Now apply FOIL to calculate the equation from these two binomials
(x-9)(x-16)=0
First x*x=x^2
Outer x*-16=-16x
Inner -9*x= -9x
Last -9*-16=144
Combine the like terms of -16x and -9x to equal -25x
Therefore, the original equation is x^2-25x+144